# Looking Backward: Retrospective Backward Synthesis for Goal-Conditioned GFlowNets

### v1

@@ Abstract
Generative Flow Networks (GFlowNets) are amortized sampling methods for learning a stochastic policy to sequentially generate compositional objects with probabilities proportional to their rewards. GFlowNets exhibit a remarkable ability to generate diverse sets of high-reward objects, in contrast to standard return maximization reinforcement learning approaches, which often converge to a single optimal solution. Recent works have arisen for learning goal-conditioned GFlowNets to acquire various useful properties, aiming to train a single GFlowNet capable of achieving different goals as the task specifies. However, training a goal-conditioned GFlowNet poses critical challenges due to extremely sparse rewards, which is further exacerbated in large state spaces. In this work, we propose a novel method named Retrospective Backward Synthesis (RBS) to address these challenges. Specifically, RBS synthesizes a new backward trajectory based on the backward policy in GFlowNets to enrich training trajectories with enhanced quality and diversity, thereby efficiently solving the sparse reward problem. Extensive empirical results show that our method improves sample efficiency by a large margin and outperforms strong baselines on various standard evaluation benchmarks.
@@ 1 Introduction
Generative Flow Networks (GFlowNets; Bengio et al. (2021)) are a new class of probabilistic models designed for sampling compositional objects from high-dimensional unnormalized distributions, and share similarities with Markov Chain Monte Carlo (MCMC) (Metropolis et al., 1953, Hastings, 1970, Andrieu et al., 2003). Unlike MCMC methods, GFlowNets do not suffer from the mixing problem (Salakhutdinov, 2009, Bengio et al., 2013, 2021), as they generate each sample independently and amortize the sampling cost. Indeed, GFlowNets transform sampling into a sequential decision-making problem: the agent learns a stochastic policy for sampling proportionally to the rewards, wherein each sequence of actions yields a unique object. In this regard, GFlowNets resemble reinforcement learning (RL), although RL agents typically seek an optimal policy maximizing the cumulative reward. There have been recent successes of GFlowNets in challenging problems due to their promising ability to generate high-quality and diverse candidates, including molecule discovery  (Bengio et al., 2021, Li et al., 2022, Jain et al., 2023), biological sequence design  (Jain et al., 2022), and causal modeling (Deleu et al., 2022, 2024, Atanackovic et al., 2024).
Humans often excel in solving compositional decision-making problems efficiently and generalizing knowledge to novel tasks due to the goal-directed nature of human behavior. This kind of goal-directed behavior is characterized by the ability to set a goal, develop a plan or strategy to achieve it and persistently take the actions that lead toward the goal. The vital role of goal-achieving has been a cornerstone not only in human development (Molinaro and Collins, 2023) but also in machine intelligence (Andrychowicz et al., 2017, Liu et al., 2022, Park et al., 2024, Niemueller et al., 2019, Addison, 2024). To pave the way towards more human-like intelligence systems for GFlowNets that exhibit improved efficiency and generalization, it is promising to incorporate goal-conditioned policies within them, which aims to achieve arbitrary goals as the task specifies.
As recently studied in OC-GAFlowNets (Pan et al., 2023a), training a goal-conditioned GFlowNet that learns to reach any goal can facilitate rapid adaptation to novel tasks with unseen rewards without the need for learning from scratch. However, it is very challenging to efficiently train GFlowNets conditioning on goals due to the sparse and binary nature of rewards, which are exacerbated in a higher-dimensional goal and state space.
@@
To tackle these critical challenges, we first formulate the problem of goal-conditioned GFlowNets (GC-GFlowNets) as learning a unified forward policy to reach any desired goals. Then, we propose a novel approach, Retrospective Backward Synthesis (RBS), which is a simple yet effective method for efficiently training GC-GFlowNets. Concretely, the insight of RBS is sourced from the data perspective, augmenting the training data with high quality and diversity. It not only transforms a sequence of actions with a negative reward into successful experiences with a positive reward, thus increasing the quality of training experiences, but also generates a totally novel trajectory to increase the diversity of data. We conduct a comprehensive evaluation of our method across various tasks from different benchmarks, where a binary bonus is given only if the agent reaches the desired goal. In comparison with a thorough set of baselines, our method largely improves sample efficiency during training and boosts the generalization ability of GFlowNets. To conclude, our contributions are three-fold:
  • We first formulate the problem of goal-conditioned GFlowNet as training universal forward policies, which take as input not only the current state but also an outcome state (goal) to achieve.
@@
  • We present a novel method named Retrospective Backward Synthesis, which imagines a new trajectory from a terminal state (goal), enhancing the quality and diversity of the training data.
@@
  • We successfully showcase the effectiveness of our method through extensive experiments. In the bit sequence generation task, our method achieves about a 100% success rate while all of the baselines completely fail. The results serve as a testament to its capability and underscore its potential for further GC-GFlowNets research.
2 Background
2.1 GFlowNets Preliminaries
@@
Let $\mathcal{X}$ denote the space of compositional objects and $R$ denote a reward function that assigns non-negative values to objects $x\in\mathcal{X}$. The non-negative reward function is denoted by $R(x)$. GFlowNets work by learning a sequential, constructive sampling policy $\pi$ that samples objects $x$ according to the distribution defined by the reward function ($\pi(x)\propto R(x)$). At each timestep, GFlowNets choose to add a building block $a\in\mathcal{A}$ (action space) to the partially constructed object $s\in\mathcal{S}$ (state space). This can be described by a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. The generation of an object $x\in\mathcal{X}$ corresponds to a complete trajectory $\tau=(s_{0}\to\cdots\to s_{n})\in\mathcal{T}$ in the DAG starting from the initial state $s_{0}$ and terminating in a terminal state $s_{n}\in\mathcal{X}$. We define state flow $F(s)$ as a non-negative weight assigned to each state $s\in\mathcal{S}$. The forward policy $P_{F}(s^{\prime}|s)$ is the forward transition probability over the children of each state, and the backward policy $P_{B}(s|s^{\prime})$ is the backward transition probability over the parents of each state. The marginal likelihood of sampling $x\in\mathcal{X}$ can be derived as $P_{F}^{\top}(x)=\sum_{\tau=(s_{0}\to\cdots\to x)}P_{F}(\tau)$. The primary objective of GFlowNets is to train a parameterized policy $P_{F}(\cdot|s,\theta)$ such that $P^{\top}_{F}(x)\propto R(x)$ (Bengio et al., 2021, 2023).
2.1.1 Training Criterion of GFlowNets
@@
Sub-Trajectory Balance. Trajectory Balance (TB) extends DB from the edge level to the trajectory level based on a telescoping calculation of Eq. (1), which parameterized the normalizing constant $Z_{\theta}$, forward policy $P_{F}(\cdot|s;\theta)$ and backward policy $P_{B}(\cdot|s;\theta)$, whose learning objective is defined as $Z_{\theta}\prod_{t=1}^{n}P_{F}(s_{t}|s_{t-1};\theta)=R(x)\prod_{t=1}^{n}P_{B}(s_{t-1}|s_{t};\theta)$. However, TB can incur large variance due to only optimizing the trajectory-level constraint (Madan et al., 2023b). Sub-trajectory Balance (SubTB) (Madan et al., 2023b) aims to mitigate the variance of TB, which considers the flow consistency criterion in the sub-trajectory level ($\tau_{i:j}=\{s_{i}\to\cdots\to s_{j}$}), where $s_{i}$ and $s_{j}$ are not necessarily the initial and terminal state. The learning objective of SubTB for each sub-trajectory is defined as in Eq. (2). It is converted into a training objective for optimization by defining the loss function as the squared difference between the left and right-hand sides of Eq. (2(Madan et al., 2023a) in the log-scale, considering a weighted combination of all possible $O(n^{2})$ sub-trajectories.
@@
3 Related Work
Generative Flow Networks (GFlowNets). Recently, there have been a number of efforts applying GFlowNets to different important cases, e.g., biological sequence design (Jain et al., 2022), molecule generation (Bengio et al., 2021), combinatorial optimization (Zhang et al., 2023a, 2024), Bayesian structure learning (Deleu et al., 2022). There have also been many works investigating how to improve the training of GFlowNets, enabling them to achieve more efficient credit assignment (Pan et al., 2023b), better exploration (Pan et al., 2023c, Lau et al., 2024), or more effective learning objectives (Bengio et al., 2023, Madan et al., 2023a) that can better handle computational complexity (Bengio et al., 2021) and large variance (Malkin et al., 2022), and generalize to stochastic environments (Pan et al., 2023d, Zhang et al., 2023b). GC-GFlowNets learn flows and policies conditioning on outcomes (goals) for reaching any targeted outcomes (Pan et al., 2023a). However, little attention has been given to this topic, leaving this promising direction largely unexplored. Meanwhile, it is challenging to train goal-conditioned policies due to sparse rewards. Our work not only provides a formal definition of GC-GFlowNets but also proposes a novel method called retrospective backward synthesis to significantly improve their training efficiency and success rates.
Goal-Conditioned Reinforcement Learning. Our formulation of goal-conditioned GFlowNets is heavily inspired by the works of goal-conditioned RL. Standard Reinforcement Learning (RL) only requires the agent to finish one specific task defined by the reward function (Schaul et al., 2015), while goal-conditioned RL trains an agent to achieve arbitrary goals as the task specifies (Andrychowicz et al., 2017). Goal-Conditioned RL augments the observation with an additional goal that the agent is required to achieve (Liu et al., 2022). The reward function is usually defined as a binary bonus of reaching the goal. To overcome the challenge of the sparsity of reward function, prior work in goal-conditioned RL has introduced algorithms based on a variety of techniques, such as hindsight relabeling (Andrychowicz et al., 2017, Fang et al., 2018, Yang et al., 2022, Fang et al., 2019, Ding et al., 2019), contrastive learning (Eysenbach et al., 2020, 2022), state-occupancy matching (Durugkar et al., 2021, Ma et al., 2022) and hierarchical sub-goal planning (Chane-Sane et al., 2021, Kim et al., 2021, Nasiriany et al., 2019). Our work is closely related to hindsight relabeling, denoted as HER (Andrychowicz et al., 2017), which relabels any experience with some commanded goal to the goal that was actually achieved in order to learn from failures. HER can generate non-negative rewards to alleviate the negative sparse reward problem, even if the agent did not complete the task. However, the agent using HER still suffers from low sample efficiency on large-scale problems, due to its limitation in only operating on the seen trajectories, while our method can imagine new trajectories with positive rewards for policy training.
4 Retrospective Backward Synthesis
In this section, we first formulate the problem of GC-GFlowNets in §4.1, followed by a motivating example in §4.2 to demonstrate the insights and efficacy of our method intuitively. Subsequently, we introduce our approach and discuss the techniques we developed for improved efficiency in §4.3.
4.1 Problem Formulation
Inspired by the literature on goal-conditioned RL  (Liu et al., 2022, Park et al., 2024, Veeriah et al., 2018), we extend the idea of flow functions and policies in GFlowNets to goal-conditioned GFowNets (GC-GFlowNets) that can generalize to different goals $y$ in the goal space (Pan et al., 2023a). To tackle this problem, we augment the DAG with goals. Formally, GC-GFlowNets are formulated as a goal-augmented DAG $\mathcal{G}=(\mathcal{S},\mathcal{A},\mathcal{Y},\phi)$, where $\mathcal{Y}$ denotes the goal space describing the tasks, and $\phi:\mathcal{S}\to\mathcal{Y}$ is a tractable mapping function that maps the state to a specific goal. In this paper, we consider an identity function for $\phi$ following Pan et al. (2023a). In the goal-augmented DAG, the reward function $R(x,y):\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is goal-conditioned, determining whether the goal object is reached:
@@ 4.1 Problem Formulation
In practice, we train GC-GFlowNets by minimizing the following loss function $\mathcal{L}$ in the log-scale shown in Eq. (5).
@@
where $F_{\theta}(s^{\prime}|y)$ is substituted with $R(s^{\prime},y)$ if $s^{\prime}$ is a terminal state.
@@ 4.2 A Motivating Example
Figure 1: Succee rates with different set sizes in set generation.
Training GC-GFlowNets according to Eq. (4) can be challenging due to the sparsity of reward signals – the agent receives a non-zero reward only when it reaches the desired goal state, while all other states yield zero reward. The high-dimensional space presents a further challenge, as the agent may spend a significant amount of time exploring unproductive regions of the state space without receiving any meaningful feedback.
Reward relabeling (Andrychowicz et al., 2017, Fang et al., 2018, Pan et al., 2023a) can alleviate sparse reward issues in goal-conditioned tasks by relabeling the achieved states as goals, but struggles to generalize to larger-scale scenarios with a large state space. This challenge is exacerbated in the context of GFlowNets, where multiple trajectories can lead to the same goal, and their number increases exponentially in dimensionality. Agents need to explore and remember all possible paths to gain comprehensive knowledge to understand and complete the task.
We demonstrate this inefficiency problem in a goal-conditioned set generation task (Pan et al., 2023b), where the agent generates a set of size $|S|$ from $|U|$ distinct elements sequentially starting from an empty set. At each timestep, the agent chooses to add an element from $U$ to the current set $s$ (the GFlowNets state) without repeating elements. We randomly sample a target state of size $|S|$ from $U$ ($|U|=30$) for each episode, and the GC-GFlowNets agent receives a negative reward of 00 as long as the final generated object is not the target state. Standard reward relabeling methods (HER (Andrychowicz et al., 2017)) struggle in the task for $|S|\geq 12$. These methods solely rely on experiences collected from interactions with environments, potentially trapping the agent in local optima and hindering the discovery of an effective goal-achieving strategy for problems with increasing scales.

### v2

@@ Abstract
Generative Flow Networks (GFlowNets), a new family of probabilistic samplers, have demonstrated remarkable capabilities to generate diverse sets of high-reward candidates, in contrast to standard return maximization approaches (e.g., reinforcement learning) which often converge to a single optimal solution. Recent works have focused on developing goal-conditioned GFlowNets, which aim to train a single GFlowNet capable of achieving different outcomes as the task specifies. However, training such models is challenging due to extremely sparse rewards, particularly in high-dimensional problems. Moreover, previous methods suffer from the limited coverage of explored trajectories during training, which presents more pronounced challenges when only offline data is available. In this work, we propose a novel method called Retrospective Backward Synthesis (RBS) to address these critical problems. Specifically, RBS synthesizes new backward trajectories in goal-conditioned GFlowNets to enrich training trajectories with enhanced quality and diversity, thereby introducing copious learnable signals for effectively tackling the sparse reward problem. Extensive empirical results show that our method improves sample efficiency by a large margin and outperforms strong baselines on various standard evaluation benchmarks. Our codes are available at https://github.com/tinnerhrhe/Goal-Conditioned-GFN.
@@ 1 Introduction
Generative Flow Networks (GFlowNets; Bengio et al. (2021)) are a new class of probabilistic models designed for sampling compositional objects from high-dimensional unnormalized distributions. GFlowNets generate each sample independently and amortize the sampling cost, and therefore do not suffer from the mixing problem (Salakhutdinov, 2009; Bengio et al., 2013; 2021) in Markov Chain Monte Carlo (MCMC) (Metropolis et al., 1953; Hastings, 1970; Andrieu et al., 2003). Indeed, GFlowNets transform sampling into a sequential decision-making problem: the agent learns a stochastic policy for sampling proportionally to the rewards, wherein each sequence of actions yields a unique object. In this regard, GFlowNets resemble reinforcement learning (RL) (Tiapkin et al., 2024; Deleu et al., 2024b; He et al., 2024), although standard RL typically focuses on optimizing policies for a single reward-maximizing objective. Due to the promising ability of GFlowNets to generate high-quality and diverse candidates, they have achieved great success in challenging problems, including molecule discovery  (Bengio et al., 2021; Li et al., 2022; Jain et al., 2023), biological sequence design  (Jain et al., 2022), and causal modeling (Deleu et al., 2022; 2024a; Atanackovic et al., 2024).
Goal-directed learning has been well conceptualized and studied across various fields, particularly in reinforcement learning (Andrychowicz et al., 2017; Park et al., 2024; Niemueller et al., 2019; Addison, 2024), where it is known as goal-conditioned RL (GCRL) (Liu et al., 2022). GCRL trains a single model and learns general policies capable of reaching arbitrary target states, showcasing significant benefits and performance improvements. However, it remains largely unexplored in the context of GFlowNets with only a few prior studies. One such study proposes training goal-conditioned GFlowNets (GC-GFlowNets) (Pan et al., 2023a) that can reach any goal within the object space (Roy et al., 2023). It further demonstrates that GC-GFlowNets facilitate rapid adaptation to novel tasks with unseen rewards, eliminating the need to learn from scratch, unlike traditional GCRL.
However, it is challenging to train GFlowNets conditioned on goals due to the sparse and binary nature of rewards, as the agent only receives positive rewards upon reaching the specified goals. GC-GFlowNets collect data by interacting with the environments in restricted steps, leading to a risk of getting trapped in constrained distributions (Yarats et al., 2021). Furthermore, previous approaches highly rely on the diversity and coverage of training trajectories, which poses a critical challenge when limited to offline data. This is particularly important in offline goal-conditioned learning, which enables training general goal-reaching policies from purely offline interaction trajectories without any further environment interaction (Ma et al., 2022a).
@@
To tackle these challenges, we propose a novel approach called Retrospective Backward Synthesis (RBS), a simple yet effective method for efficiently training GC-GFlowNets that learns a unified forward policy capable of reaching any desired goals. Different from existing approaches inspired by GCRL literature (Pan et al., 2023a; Roy et al., 2023), RBS augments the forward trajectories collected by the forward policies with synthesized backward trajectories by looking backward, guided by the inherent backward policies. It is noteworthy that the synthesized backward trajectories represent successful experiences as they consistently reach the desired goals. The key insight of RBS lies in its data-driven approach, which enriches the training data with high quality and diversity. It not only transforms sequences of actions with failure rewards into successful experiences with meaningful rewards, thereby increasing the quality of training experiences, but also generates entirely novel trajectories to increase the diversity of data. Nevertheless, it is still challenging to scale it up to more complex and long-horizon problems, where GC-GFlowNets are prone to instabilities and mode collapse. We hypothesize this is due to ineffective reward gradient propagation caused by severe credit assignment issues, which results in poor utilization of valuable learning signals. To address these issues, we propose to intensify reward signals in the training objective to strengthen gradient backpropagation and regularize the backward policies to enhance the diversity of synthesized backward trajectories. We conduct a comprehensive evaluation of our method across various tasks from different benchmarks, where a binary bonus is awarded only if the agent reaches the desired goal. In comparison with a thorough set of baselines, our method largely improves sample efficiency during training and enhances the generalizability of GC-GFlowNets. Our contributions are three-fold:
@@
  • We present a novel method named Retrospective Backward Synthesis, which imagines a new trajectory from a desired goal, enhancing the quality and diversity of the training data.
  • We introduce effective techniques, e.g., reward intensification and backward policy regularization, to stabilize and improve the training process.
@@
  • We showcase the effectiveness of our method through extensive experiments. A noteworthy result is that our method achieves about a 100% success rate in large-scale sequence generation tasks while all of the baselines completely fail. The results serve as a testament to its capability and underscore its potential for further GC-GFlowNets research.
2 Preliminaries
2.1 GFlowNets
@@
Let $\mathcal{X}$ denote the space of compositional objects and $R$ denote a reward function that assigns non-negative values to objects $x\in\mathcal{X}$. The non-negative reward function is denoted by $R(x)$. GFlowNets work by learning a sequential, constructive sampling policy $\pi$ that samples objects $x$ according to the distribution defined by the reward function ($\pi(x)\propto R(x)$). At each timestep, GFlowNets choose to add a building block $a\in\mathcal{A}$ (action space) to the partially constructed object $s\in\mathcal{S}$ (state space). This can be described by a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $\mathcal{S}$ is a finite set of all possible states, and $\mathcal{A}$ is a subset of $\mathcal{S}\times\mathcal{S}$, representing directed edges. The generation of an object $x\in\mathcal{X}$ corresponds to a complete trajectory $\tau=(s_{0}\to\cdots\to s_{n})\in\mathcal{T}$ in the DAG starting from the initial state $s_{0}$ and terminating in a terminal state $s_{n}\in\mathcal{X}$. We define state flow $F(s)$ as a non-negative weight assigned to each state $s\in\mathcal{S}$. The forward policy $P_{F}(s^{\prime}|s)$ is the forward transition probability over the children of each state, and the backward policy $P_{B}(s|s^{\prime})$ is the backward transition probability over the parents of each state. The marginal likelihood of sampling $x\in\mathcal{X}$ can be derived as $P_{F}^{\top}(x)=\sum_{\tau=(s_{0}\to\cdots\to x)}P_{F}(\tau)$. The primary objective of GFlowNets is to train a parameterized policy $P_{F}(\cdot|s,\theta)$ such that $P^{\top}_{F}(x)\propto R(x)$ (Bengio et al., 2021; 2023).
2.1.1 Training Criteria of GFlowNets
@@
(Sub-) Trajectory Balance. Trajectory Balance (TB) extends DB from the edge level to the trajectory level based on a telescoping calculation of Eq. (1), which parameterized the normalizing constant $Z_{\theta}$, forward policy $P_{F}(\cdot|s;\theta)$ and backward policy $P_{B}(\cdot|s;\theta)$, whose learning objective is defined as $Z_{\theta}\prod_{t=1}^{n}P_{F}(s_{t}|s_{t-1};\theta)=R(x)\prod_{t=1}^{n}P_{B}(s_{t-1}|s_{t};\theta)$. Sub-trajectory Balance (SubTB) (Madan et al., 2023b) aims to mitigate the variance of TB and bias of DB, which considers the flow consistency criterion in the sub-trajectory level ($\tau_{i:j}=\{s_{i}\to\cdots\to s_{j}$}), where $s_{i}$ and $s_{j}$ are not necessarily the initial and terminal state. The learning objective of SubTB for each sub-trajectory is defined as in Eq. (2). Its loss function is the squared difference between the left and right-hand sides of Eq. (2(Madan et al., 2023a) in the log-scale, considering a weighted combination of all possible $O(n^{2})$ sub-trajectories.
@@
2.2 Problem Formulation of GC-GFlowNets
Inspired by the literature on goal-conditioned RL  (Liu et al., 2022; Park et al., 2024; Veeriah et al., 2018), the idea of flow functions and policies in GFlowNets can be generalized to different goals $y$ in the goal space (Pan et al., 2023a), leading to the formulation of goal-conditioned GFowNets (GC-GFlowNets). GC-GFlowNets can be formulated as a goal-augmented DAG $\mathcal{G}=(\mathcal{S},\mathcal{A},\mathcal{Y},\phi)$, where $\mathcal{Y}$ denotes the goal space describing the tasks, and $\phi:\mathcal{S}\to\mathcal{Y}$ is a tractable mapping function that maps the state to a specific goal. In this paper, we consider an identity function for $\phi$ following Pan et al. (2023a). In the goal-augmented DAG, the reward function $R(x,y):\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is also conditioned on goals, determining whether the goal object is reached:
@@ 2.2 Problem Formulation of GC-GFlowNets
In practice, GC-GFlowNets can be trained by minimizing the following loss function $\mathcal{L}$ in the log-scale (for numerical stability as discussed in Bengio et al. (2021)) as shown in Eq. (5), where $F_{\theta}(s^{\prime}|y)$ is substituted with $R(s^{\prime},y)$ if $s^{\prime}$ is a terminal state.
@@
3 Proposed Method: Retrospective Backward Synthesis
In this section, we first introduce a motivating example in §3.1 to demonstrate the insights and efficacy of our method intuitively. Subsequently, we introduce our novel approach and discuss the techniques we developed for improved efficiency in §3.2, which improves the training of GC-GFlowNets in a simple yet effective manner.
@@ 3.1 A Motivating Example
Training GC-GFlowNets according to Eq. (4) can be challenging due to the sparsity of reward signals – the agent receives a non-zero reward only when it reaches the desired goal state, while all other states yield zero reward. The high-dimensional space presents a further challenge, as the agent may spend a significant amount of time exploring unproductive, restricted regions of the state space without receiving any meaningful feedback. Reward relabeling (Andrychowicz et al., 2017; Fang et al., 2018; Pan et al., 2023a) aims to alleviate sparse reward issues in goal-conditioned tasks by relabeling the achieved states as goals, which have demonstrated success in the goal-conditioned RL literature, but still struggles to generalize to larger-scale scenarios with a large state space. This challenge is exacerbated in the context of GFlowNets, where agents can discover different trajectories leading to the same goal, and their number increases exponentially in dimensionality. Consequently, these reward labeling techniques may struggle to generalize across these different goal-trajectory relationships and are prone to get stuck in local optima, as they rely solely on observed data without expanding their data coverage.
Figure 1: Succee rates with increasing set sizes in set generation.
We demonstrate this inefficiency problem in a goal-conditioned set generation task (Pan et al., 2023b), where the agent generates a set of size $|S|$ from $|U|$ distinct elements sequentially starting from an empty set. At each timestep, the agent chooses to add an element from $U$ to the current set $s$ (the GFlowNets state) without repeating elements. We randomly sample a target state of size $|S|$ from $U$ ($|U|=30$) for each episode, and the GC-GFlowNets agent receives a negative reward of 00 as long as the final generated object is not the target state. Previous state-of-the-art method (Pan et al., 2023a) based on standard reward relabeling (HER (Andrychowicz et al., 2017)), struggles in this high-dimensional task for $|S|\geq 12$ with large state space. These methods solely rely on experiences collected from interactions with environments, potentially trapping the agent in local optima and hindering the discovery of an effective goal-achieving strategy for problems with increasing scales.